Elasticity A1.25m aluminum rol (Y = 7.0 × 1011) that is 3 crn in diameter has a large cement block with a mass of 10,000 kg balanced on its end compressing the rod. What are the stress and strain experienced by the block? By what amount is the rod compressed?
Answer #1

σ = Fn / A
Fn = 10000 kg (9.8 m/s2) = 98000 N
A = pi(r^2) = 3.14*0.015*0.015 = 0.0007685 m2
σ = 98000 N/0.0007685 m2 = 138641639.315 N/m2

Strain, ε = dl / lo

= σ / Y = 138641639.315 N/m2/7.0*10^11 = 1.98*10^-4

dl = compression =  2.47*10^-4 m = 0.247 mm